OpenStudy (anonymous):

What is the distance between the points P(-2, 3) and Q (-5, -6)?

OpenStudy (anonymous):

sqrt((y2-y1)^2 + (x2-x1)^2) The 2 points are endpoints of a line. This line forms the hypotenuse of a right triangle. That hypotenuse is \[\sqrt{ \Delta y ^{2} + \Delta x ^{2} }\]

OpenStudy (anonymous):

to get the distance between to (x,y) co-ordinates you can use the distance formula: \[\sqrt{(y _{2}-y _1)^2 + (x_2-x_1)^2}\] then you input your values like so: \[\sqrt{(-6-3)^2 + (-5-(-2))^2}=......\] off you go!